OpenStudy (anonymous):

a spaceship engine slowly increased the velocity by about +90m/s per day (a) how much time in days would it take to increase the velocity of the spaceship by +2700m/s (b) what was the acceleration of the spaceship in m/s squared

OpenStudy (mani_jha):

time taken = 2700/90 = 30 days = 30*86400 seconds acceleration = 90/86400